# Survey on the Evaluation of Generative Models in Music

### v1

@@ 4. Evaluation of usability & user experience
In this section we describe Human-Computer Interaction (HCI) methods and techniques that have been used to evaluate user interaction with generative music models. These methods and techniques share aims with Human-Centered AI (Shneiderman, 2022; Ozmen Garibay et al., 2023) to research, design, and evaluate AI systems from a human-centered or user-first perspective. Below, we first introduce HCI methodologies for evaluating AI music systems. We then describe data collection methods for understanding the user experience of generative AI including quantitative and qualitative approaches. It is important to note that there are no de-facto standards defining which data collection methods are used for which evaluation methodologies. Instead, selection of data collection and methodology is based on best practice in the field and individual HCI practitioner skills and experience.
@@ 4.1.2. Online evaluation
Online evaluation settings can be helpful for evaluating generative AI systems across a large sample size. For example, Ben-Tal et al. (2021) examined a version of the FolkRNN generative AI model hosted online, where they were able to examine how people generated content with FolkRNN serendipitously and how they tweaked values to modify and curate outputs. Typically, generative AI systems are either hosted online for user interaction or available as a download to users. Audiences across the globe can then be reached using survey platforms such as Prolific, [^2] https://www.prolific.com/, last accessed: Jun 25, 2024 and filtered for characteristics such as nationality or technical skills. Whilst online settings allow for large numbers of participants, they lack the rigor of controlled experiments.
@@ 4.1.3. Exploratory studies
Exploratory studies emphasize evaluating user experience and collecting subjective feedback. They typically involve more open-ended tasks than controlled experiments or online evaluations. Exploratory studies could, for example, take place in a controlled lab setting, yet be more open-ended to allow music making to occur in a more natural way instead of comparing two conditions directly, or may take place in less controlled setting. For example, Bougueng Tchemeube et al. (2023) tasked people with exploring their generative AI interface and collected structured questionnaire measures to capture aspects of people’s user experience, but do not make strict comparisons between interface designs. Note that a balance can be struck between controlled and exploratory evaluations, such as giving open-ended tasks in controlled study settings to test generative AI in a way that is closer to real-world music making (e.g., (Louie et al., 2020)), or to give structure to data collection in real-world settings (e.g., (Ford et al., 2024)).
@@
Research-in-the-wild (Benford et al., 2013; Chamberlain et al., 2012) approaches contrast controlled experiments to evaluate generative AI models in their real-world places of use, possibly over extended periods of time. For example, in ethnographic approaches (Chamberlain et al., 2012; Benford et al., 2013) the researcher takes observations or field notes, or collects data on patterns of behavior that people have naturally exhibited while making music. For AI music, this type of approach has been used in, e.g., the context of the international AI music song contest (Huang et al., 2020). Here, the researchers were able to identify ways that developers and musicians worked together to create music, for example by preferring to curate AI generated content instead of (re-)developing their AI tools. Across the HCI studies on generative music, there are several examples of ethnographic-inspired observations being collected. Fiebrink et al. (2010), for instance, “recorded text minutes of [composer’s] activities, discussion topics, and specific questions, problem reports, and feature requests” for seminars on their Wekinator (Fiebrink et al., 2011) system. Bryan-Kinns et al. (2024b) used first-person accounts of music making with a generative AI system over several months to understand how it was appropriated into music making practice.
@@ 4.2. User data collection: Quantitative
The key quantitative data collection technique used to evaluate generative AI systems are questionnaires. These are used to quantify both subjective feelings of a system’s usability as well as more experiential aspects. Typically, these questionnaires place statements against a Likert-type scale (Likert, 1932) that rates on a scale, e.g., from 1–5, how much a user agrees or disagrees with a statement.
@@
User experience-oriented scales include the User Experience Questionnaire (Laugwitz et al., 2008) and the User Engagement Scale (O’Brien et al., 2018). The most influential questionnaire with respect to creative technology within the last 10 years is arguably the Creativity Support Index (CSI) (Cherry and Latulipe, 2014): a questionnaire designed to test the capacity of a tool to support creativity, offering factors for several important aspects of the creative user experience including: focused attention, enjoyment and collaboration. For AI music, however, we surprisingly find few examples of AI music user studies that have adopted the CSI — Bougueng Tchemeube et al. (2023) being a notable exception. We instead observe many examples where researchers have chosen to define their own questionnaires to explore constructs for the AI systems that are not captured in current standardized scales. For instance, Louie et al. (2020) invented their own questions to capture a person’s feelings of agency. Ford and Bryan-Kinns (2023) identified reflection as an aspect missing from the CSI whilst being an important factor in AI music making (Ford et al., 2024; Ford and Bryan-Kinns, 2022). Alongside established measures of creativity support and usability, Bougueng Tchemeube et al. (2023) also added questions central to human-AI interaction on trust, perceived authorship, and flexibility. Questionnaires are more often used in usability-focused evaluation methodologies such as controlled experiments and less frequently used as part of more experiential evaluations such as in-the-wild studies. HCI researchers also use questionnaires to establish characteristics of their sample under study. For example, the Goldsmith’s Musical Sophistication Index (Müllensiefen et al., 2014) offers a standardized metric for musical expertise, helping to identify whether the users of a generative AI system under study have above or below average musical skills.
@@ 4.3. User data collection: Qualitative
In addition to quantitative data collection researchers use qualitative data collection to gain greater insight into users’ feelings, motivations, and reflections when using generative AI system. It is important to note that triangulation across different data collection approaches (also referred to as mixed methods) is crucial — using both qualitative and quantitative data can help to demonstrate which features of the user experience are or are not improved by AI, as well as offering insights into why this might be so (Bryan-Kinns and Reed, 2023).
@@ 4.3.3. First-person perspectives
The use of methods such as questionnaires and interviews described above are borne from a psychology-driven epistemological stance: to identify generalizable models of how people interact with technology. Approaches inspired by Arts and Humanities offer insights into the individuality and subjectivity of how artists have interpreted their use of technology (Candy and Edmonds, 2018). We identified an increasing trend to report on the use of AI from a first-person perspective (Noel-Hirst and Bryan-Kinns, 2023; Ford et al., 2024; Sturm et al., 2019a; Sturm, 2022; Ben-Tal et al., 2021; Bryan-Kinns et al., 2024b), publishing perspectives on how individuals have been able to use and incorporate models into their music-making. In some cases, these are autoethnographies (Noel-Hirst and Bryan-Kinns, 2023; Sturm, 2022) where a researcher reflects on their own practice by means of capturing data over a long time period. Other examples show collections of first-person accounts (Ford et al., 2024; Sturm et al., 2019a). Sturm et al. (2024)’s proposal for a field of AI music studies engages with these methods to explore ways to more meaningfully and critically engage with the broader communities in social sciences and humanities. In contrast to questionnaires, first-person perspectives are more frequently used in user experience-focused evaluations such as in-the-wild studies and less frequently, if at all, in controlled experiments.
@@ 4.4. Other HCI evaluation approaches
There is a wide variety of other evaluation methods used in HCI beyond those we found are most often used when evaluating AI music systems as described in this section. Usability metrics such as task completion rates, task time or the number of errors, have been explored to evaluate the usability of a computer music system (Wanderley and Orio, 2002). However, these are not prominent in generative AI music user studies where music making tends to have no clearly defined goal. “Wizard of Oz” was an early popular approach to user studies of human-AI interaction more broadly, where users interact with a user interface whilst a researcher provides feedback through the interface in lieu of an AI. For example, Thelle and Fiebrink (2022) tested participant’s reactions to researchers who performed piano phrases live, acting as an AI system. Similar examples have been tested with more complex generative music programming languages (Bellingham, 2022). We also did not find many examples of using physiological measures such as heart rate or eye-tracking to explore people’s interaction with music AI tools. In other arts-based HCI studies, physiological measures have been used as proxies of aspects of the user experience such as anxiety or boredom, indicated by participants’ heart rate or skin conductance (Maier et al., 2019). This could be an open area for further research.
@@ 5.1. Validity
Although subjective evaluation is often considered the most meaningful way of evaluating the output of generative systems, the results cannot be automatically assumed to be robust or reliable. Subjective evaluation of system output (Sect. 3.1) usually relies on survey approaches (Hernandez-Olivan et al., 2022), much as quantitative user data collection relies on questionnaires (Sect. 4.2). However, survey creation and questionnaire design is non-trivial — the creation of batteries and psychometrics in psychology is an entire subfield (Gehlbach and Brinkworth, 2011; Sanbonmatsu et al., 2021). Moreover, questionnaires have not yet been developed to cover many experiential aspects of human-AI interaction such as people’s feelings of agency. A considerable number of researchers working on generative music systems lack the background, skill and/or resources to successfully carry out such surveys with valid, reliable, and replicable results (Yang and Lerch, 2020).
In addition, there is a known bias in people’s perception against AI-generated music (Shank et al., 2023). As such, any test in which a question about human vs. machine creation is raised may introduce bias in the results. With respect to Turing tests, Hernández-Orallo (2020) point out that a known validity issue with Turing tests is that the outcomes cannot disambiguate whether the model was a good imitator, or the human was a poor judge.
Given these potential validity issues with subjective evaluation of system output, objective evaluation remains a viable choice to complement listening studies. With objective evaluation, however, there are other validity concerns. Some of these concerns start with the data and its characteristics: is the sample size large enough, does the data reflect the targeted homogeneity or heterogeneity, are there confounding characteristics in the data that complicate drawing conclusions, etc. Another concern is how meaningful the used metrics are, i.e., whether the objective metrics properly represent the evaluation target or whether differences in evaluation results are perceptually relevant or not. Even if some metrics prove to be relevant, individual metrics or criteria are usually insufficient to provide a broader assessment; Theis et al. (2016) rightly note that “Good performance with respect to one criterion (…) need not imply good performance with respect to the other criteria.” Note that even when targeting very specific criteria (e.g., complexity), subjective impressions might be better predictors of responses than objective measures (Hargreaves, 2010). For that reason, it might make more sense to use subjective impressions of criteria (rated by the listener) as predictors of the overall judgments of aesthetic value (Juslin et al., 2023).
@@ 5.2. Perceptual and musical relevance of objective metrics
Revisiting the objective metrics introduced in Sect. 3, it can be observed that the most common metrics compare training data characteristics with characteristics of the generated data in one way or the other. A major differentiation between such metrics is the space and the dimensions in which different metrics approach such a comparison. On the one hand, there exist learned embeddings such as VGGish (Hershey et al., 2017) for the FAD, on the other hand there are custom-designed low-level statistical descriptors such as pitch range, pitch class histograms, etc. (Yang and Lerch, 2020). This has considerable impact on the interpretability of the descriptors; while a learned descriptor such as VGGish cannot be interpreted directly, custom-designed descriptors tend to be more interpretable. However, high interpretability does neither mean that the descriptor is relevant for assessment nor that it is perceptually or musically meaningful.
Perceptual studies of specific descriptors are necessary to understand their relevance and meaning. Simply finding a difference between two set of data with respect to one descriptor does not automatically mean that these data are different from a perceptual point of view.
Furthermore, even descriptors known for their perceptual validity can be more or less meaningful depending on context and scenario. For instance, it has been demonstrated that people are incredibly sensitive to the statistical distribution of pitches in a piece of music, and can even learn new musical systems and grammars based on such statistical inference (e.g., (Krumhansl, 2000; Huron, 2006; Temperley, 2004; Loui et al., 2023). However, in general, not all musical representations are equal, and this can have a sizeable impact in the perceptual relevance of any given feature or metric. For instance, most music is tonal meaning that it is (at least temporarily) in a given key or mode and has a stable tonic. For such tonal music, listeners are very sensitive to notes that are outside of the key (i.e., ’wrong notes’) (Krumhansl, 2000; Dowling, 1991; Raman et al., 2016). As such, measuring the statistical distribution of pitches in relation to that tonic (as scale degrees or musical intervals from a tonic) carries a different musical and perceptual relevance compared to the distribution of all pitch classes measured in a tonic-agnostic way. Thus, it is not only necessary to validate whether certain descriptors have perceptual relevance per se, but also in what (e.g., tonal or stylistic) context they are extracted. But even given a set of relevant descriptors we can only guess how exhaustive this set is. At the very least, the number and type of relevant descriptors is genre-dependent, and is quite possibly indefinite.

### v2

@@ 4. Evaluation of usability & user experience
In this section we describe Human-Computer Interaction (HCI) methods and techniques that have been used to evaluate user interaction with generative music models. These methods and techniques share aims with Human-Centered AI (Shneiderman, 2022; Ozmen Garibay et al., 2023) to research, design, and evaluate AI systems from a human-centered or user-first perspective. Below, we first introduce HCI methodologies for evaluating AI music systems. We then describe data collection methods for understanding the user experience of generative AI including quantitative and qualitative approaches. Table 3 closes this section by summarizing key HCI evaluation approaches discussed. It is important to note that there are no de-facto standards defining which data collection methods are used for which evaluation methodologies. Instead, selection of data collection and methodology is based on best practice in the field and individual HCI practitioner’s skills and experience.
@@ 4.1.2. Online evaluation
Online evaluation settings can be helpful for evaluating generative AI systems across a large sample size. For example, Ben-Tal et al. (2021) evaluated a version of the FolkRNN generative AI model hosted online, where they were able to examine how people generated content with FolkRNN serendipitously and how they tweaked values to modify and curate outputs. Typically, generative AI systems are either hosted online for user interaction or available as a download to users. Audiences across the globe can then be reached using survey platforms such as Prolific, [^2] https://www.prolific.com/, last accessed: Jun 25, 2024 and filtered for characteristics such as nationality or technical skills. Whilst online settings allow for large numbers of participants, they lack the rigor of controlled experiments.
@@ 4.1.3. Exploratory studies
Exploratory studies emphasize evaluating user experience and collecting subjective feedback. They typically involve more open-ended tasks than controlled experiments or online evaluations. Exploratory studies can, for example, take place in a controlled lab setting which is more typical of comparative studies, yet use open-ended tasks that allow music making to occur in a more natural way. For example, Bougueng Tchemeube et al. (2023) tasked people with exploring their generative AI interface and collected structured questionnaire measures to capture aspects of people’s user experience, but do not make strict comparisons between interface designs. Thus, balance can be struck between controlled and exploratory evaluations, such as giving open-ended tasks in controlled study settings to test generative AI in a way that is closer to real-world music making (e.g., (Louie et al., 2020)), or to give structure to data collection in real-world settings (e.g., (Ford et al., 2024)).
@@
Research-in-the-wild (Benford et al., 2013; Chamberlain et al., 2012) approaches contrast controlled experiments to evaluate generative AI models in their real-world places of use, possibly over extended periods of time. For example, in ethnographic approaches (Chamberlain et al., 2012; Benford et al., 2013) the researcher takes observations or field notes, or collects data on patterns of behavior that people have naturally exhibited while making music. For AI music, this type of approach has been used in, e.g., the international AI music song contest (Huang et al., 2020). The researchers identified how developers and musicians collaborated to create music , for example by preferring to curate AI generated content instead of (re-)developing their AI tools. Across the HCI studies on generative music, there are several examples of ethnographic-inspired observations being collected. Fiebrink et al. (2010), for instance, “recorded text minutes of [composer’s] activities, discussion topics, and specific questions, problem reports, and feature requests” for seminars on their Wekinator (Fiebrink et al., 2011) system. Bryan-Kinns et al. (2024b) used first-person accounts of music making with a generative AI system over several months to understand how it was appropriated into music making practice.
@@ 4.2. User data collection: Quantitative
The primary quantitative method for evaluating generative AI systems is the use of questionnaires. These are used to quantify both subjective feelings of a system’s usability as well as more experiential aspects. Typically, these questionnaires use a Likert-type scale (Likert, 1932) to measure user agreement with statements on a scale (e.g., 1–5).
@@
User experience-oriented scales include the User Experience Questionnaire (Laugwitz et al., 2008) and the User Engagement Scale (O’Brien et al., 2018). The most influential questionnaire with respect to creative technology within the last 10 years is arguably the Creativity Support Index (CSI) (Cherry and Latulipe, 2014): a questionnaire designed to test the capacity of a tool to support creativity, offering factors for several important aspects of the creative user experience including: focused attention, enjoyment and collaboration. For AI music, however, we found surprisingly few examples of AI music user studies that have adopted the CSI — Bougueng Tchemeube et al. (2023) a notable exception. We instead observe many examples where researchers have chosen to define their own questionnaires to explore constructs for the AI systems that are not captured in current standardized scales. For instance, Louie et al. (2020) invented their own questions to capture a person’s feelings of agency. Ford and Bryan-Kinns (2023) identified reflection as an aspect missing from the CSI whilst being an important factor in AI music making (Ford et al., 2024; Ford and Bryan-Kinns, 2022). Alongside established measures of creativity support and usability, Bougueng Tchemeube et al. (2023) also added questions central to human-AI interaction on trust, perceived authorship, and flexibility. This use of researcher defined questionnaires reflects the dominance of measures of engagement and usability in assessing AI interaction in creativity-related HCI research (Cox et al., 2025; Rhys Cox et al., 2025).
Questionnaires are more often used in usability-focused evaluation methodologies such as controlled experiments and less frequently used as part of more experiential evaluations such as in-the-wild studies. HCI researchers also use questionnaires to establish characteristics of their sample under study. For example, the Goldsmith’s Musical Sophistication Index (Müllensiefen et al., 2014) offers a standardized metric for musical expertise, helping to identify whether the users of a generative AI system under study have above or below average musical skills.
@@ 4.3. User data collection: Qualitative
In addition to quantitative data collection researchers use qualitative data collection to gain greater insight into users’ feelings, motivations, and reflections when using a generative AI system. It is important to note that triangulation across different data collection approaches (also referred to as mixed methods) is crucial — using both qualitative and quantitative data can help to demonstrate which features of the user experience are improved by AI or not, as well as offering insights into why this might be so (Bryan-Kinns and Reed, 2023).
@@ 4.3.3. First-person perspectives
The use of methods such as questionnaires and interviews described above is borne from a psychology-driven epistemological stance: to identify generalizable models of how people interact with technology. Approaches inspired by Arts and Humanities offer insights into the individuality and subjectivity of how artists have interpreted their use of technology (Candy and Edmonds, 2018). We identified an increasing trend to report on the use of AI from a first-person perspective (Noel-Hirst and Bryan-Kinns, 2023; Ford et al., 2024; Sturm et al., 2019a; Sturm, 2022; Ben-Tal et al., 2021; Bryan-Kinns et al., 2024b), publishing perspectives on how individuals have been able to use and incorporate models into their music-making. In some cases, these are autoethnographies (Noel-Hirst and Bryan-Kinns, 2023; Sturm, 2022) where a researcher reflects on their own practice by means of capturing data over a long time period. Other examples show collections of first-person accounts (Ford et al., 2024; Sturm et al., 2019a). Sturm et al. (2024)’s proposal for a field of AI music studies engages with these methods to explore ways to more meaningfully and critically engage with the broader communities in social sciences and humanities. In contrast to questionnaires, first-person perspectives are more frequently used in user experience-focused evaluations such as in-the-wild studies and less frequently, if at all, in controlled experiments.
Table 3. Overview of Human-Computer Interaction evaluation approaches.
Quantitative Methods: Questionnaires
Acronym Captures Reference
Goldsmith’s Musical Sophistication Index GMSI Musical expertise (Müllensiefen et al., 2014)
NASA Task Load Index NASA TLX How complex a task is perceived to be (Hart, 2006)
Standard System Usability Scale SUS The usability of a user interface (Bangor et al., 2008)
Cognitive Dimensions of Notations questionnaire CD The usability of a user interface (Blackwell and Green, 2000)
User Experience Questionnaire UEQ The usability and experiential aspects of a user interface (Laugwitz et al., 2008)
User Engagement Scale UES The experiential aspects of a user interface (O’Brien et al., 2018)
Creativity Support Index CSI How well a user interface supports creative work (Cherry and Latulipe, 2014)
Reflection in Creative Experience RiCE Types of reflection in creative contexts (Ford and Bryan-Kinns, 2023)
Qualitative Methods
Examples Captures Reference
Interviews: Structured, semi-structured, or open-ended (Suh et al., 2021; Fiebrink and Sonami, 2020; Louie et al., 2020; Frid et al., 2020; Ford et al., 2024) Insights into users’ thoughts, motivations, and feelings on their interaction (Braun and Clarke, 2013)
Think-aloud (Suh et al., 2021; Louie et al., 2020, 2022; Frid et al., 2020; Hunt, 2021) Users’ thought processes whilst using a user interface (Preece et al., 2015)
Video-cued recall (Candy et al., 2006) Users’ post-hoc thoughts about using a user interface (Alshammari et al., 2015)
Autoethnographies (Noel-Hirst and Bryan-Kinns, 2023; Sturm, 2022) A researcher’s subjective and personal reflections on their musical practice and use of technology (Rapp, 2018)
First-person accounts (Ford et al., 2024; Sturm et al., 2019a) Rich descriptions of users’ personal reflections on using AI models (Lucero et al., 2019)
@@ 4.4. Other HCI evaluation approaches
There is a wide variety of other evaluation methods used in HCI beyond those we found are most often used when evaluating AI music systems, as described in this section. Usability metrics such as task completion rates, task time or the number of errors, have been explored to evaluate the usability of a computer music system (Wanderley and Orio, 2002). However, these are not prominent in generative AI music user studies where music making tends to have no clearly defined goal. “Wizard of Oz” was an early popular approach to user studies of human-AI interaction more broadly, where users interact with a user interface whilst a researcher provides feedback through the interface in lieu of an AI. For example, Thelle and Fiebrink (2022) tested participant’s reactions to researchers who performed piano phrases live, acting as an AI system. Similar examples have been tested with more complex generative music programming languages (Bellingham, 2022). We also did not find many examples of using physiological measures such as heart rate or eye-tracking to explore people’s interaction with music AI tools. In other arts-based HCI studies, physiological measures have been used as proxies of aspects of the user experience such as anxiety or boredom, indicated by participants’ heart rate or skin conductance (Maier et al., 2019). This could be an open area for further research.
@@ 5.1. Validity
Although subjective evaluation is often considered the most meaningful way of evaluating the output of generative systems, the results cannot be automatically assumed to be robust or reliable. Subjective evaluation of system output (Sect. 3.1) usually relies on survey approaches (Hernandez-Olivan et al., 2022), much as quantitative user data collection relies on questionnaires (Sect. 4.2). However, designing surveys and questionnaires is non-trivial — the creation of batteries and psychometrics form an entire subfield of psychology (Gehlbach and Brinkworth, 2011; Sanbonmatsu et al., 2021). Moreover, existing questionnaires often fail to capture experiential aspects of human-AI interaction, such as users’ sense of agency. A considerable number of researchers working on generative music systems lack the background, skill and/or resources to successfully carry out such surveys with valid, reliable, and replicable results (Yang and Lerch, 2020).
In addition, there is a known bias in people’s perception against AI-generated music (Shank et al., 2023). As such, tests highlighting human vs. machine authorship may introduce bias in the results. With respect to Turing tests, Hernández-Orallo (2020) point out that a known validity issue with Turing tests is that the outcomes cannot disambiguate whether the model was a good imitator, or the human was a poor judge.
Given these potential validity issues with subjective evaluation of system output, objective evaluation remains a viable choice to complement listening studies. With objective evaluation, however, there are other validity concerns. These concerns often start with the data and its characteristics: Is the sample size sufficient? Do the data reflect the targeted homogeneity or heterogeneity? Are there confounding characteristics in the data that complicate drawing conclusions? Another concern is the validity of the chosen metrics — do they meaningfully represent the evaluation target, and are observed differences in evaluation results perceptually significant? Even if some metrics prove to be relevant, individual metrics or criteria provide sufficient breadth for comprehensive evaluation; Theis et al. (2016) rightly note that “Good performance with respect to one criterion (…) need not imply good performance with respect to the other criteria.” Note that even when targeting very specific criteria (e.g., complexity), subjective impressions might be better predictors of responses than objective measures (Hargreaves, 2010). For that reason, it might make more sense to use subjective impressions of criteria (rated by the listener) as predictors of the overall judgments of aesthetic value (Juslin et al., 2023).
@@ 5.2. Perceptual and musical relevance of objective metrics
Revisiting the objective metrics introduced in Sect. 3, it can be observed that the most common metrics compare training data characteristics with characteristics of the generated data in one way or the other. A major differentiation between such metrics is the space and the dimensions in which different metrics approach such a comparison. On the one hand, learned embeddings such as VGGish (Hershey et al., 2017) and a variety of other embeddings have been utilized for the FAD or related metrics, on the other hand there are custom-designed low-level statistical descriptors such as pitch range, pitch class histograms, etc. (Yang and Lerch, 2020). This has considerable impact on the interpretability of the descriptors; while a learned descriptor such as VGGish cannot be interpreted directly, custom-designed descriptors tend to be more interpretable. However, high interpretability does neither mean that the descriptor is relevant for assessment nor that it is perceptually or musically meaningful.
Perceptual studies of specific descriptors are necessary to understand their relevance and meaning. Simply finding a difference between two set of data with respect to one descriptor does not automatically mean that these data are different from a perceptual point of view. Recent studies on the suitability of learned embeddings for evaluation seem to exclusively focus on overall relevance for summary aesthetic judgments without considering interpretability or musical meaning (Huang et al., 2025; Grötschla et al., 2025).
Furthermore, even descriptors known for their perceptual validity can be more or less meaningful depending on context and scenario. For instance, it has been demonstrated that people are incredibly sensitive to the statistical distribution of pitches in a piece of music, and can even learn new musical systems and grammars based on such statistical inference (e.g., (Krumhansl, 2000; Huron, 2006; Temperley, 2004; Loui et al., 2023)). However, in general, not all musical representations are equal, and this can have a sizeable impact in the perceptual relevance of any given feature or metric. For instance, most music is tonal meaning that it is (at least temporarily) in a given key or mode and has a stable tonic. For such tonal music, listeners are very sensitive to notes that are outside of the key (i.e., ’wrong notes’) (Krumhansl, 2000; Dowling, 1991; Raman et al., 2016). As such, measuring the statistical distribution of pitches in relation to that tonic (as scale degrees or musical intervals from a tonic) carries a different musical and perceptual relevance compared to the distribution of all pitch classes measured in a tonic-agnostic way. Thus, it is not only necessary to validate whether certain descriptors have perceptual relevance per se, but also in what (e.g., tonal or stylistic) context they are extracted. But even given a set of relevant descriptors we can only guess how exhaustive this set is. At the very least, the number and type of relevant descriptors is genre-dependent, and is quite possibly indefinite.